# CaMMT: Benchmarking Culturally Aware Multimodal Machine Translation

### v1

@@ 7 Conclusions
We present CaMMT, a human-curated dataset for Multimodal Machine Translation that encompasses 19 languages across 23 regions. We evaluated five VLMs at different scales on CaMMT and observed that providing images as auxiliary context generally improves translation quality in ways that native speakers find meaningful. When translations incorporate visual context, they tend to better preserve cultural elements, use correct gender forms, and resolve ambiguities—improvements that automatic MT metrics often miss. However, we also observe a non-trivial number of cases where visual input negatively affects translations. Understanding when and why this occurs remains an important direction for future research.
@@
Appendix A CaMMT Statistics
@@
Language-Region Script(s) Size CSI Culturally Relevant (non-CSI) Not culturally relevant
@@ Appendix B Experimental Setting
Appendix C BLEU and BertScore metrics across models
@@ Appendix D Translation Prompts
Appendix E Comparison between categories measured by chrF scores
@@
Appendix F License
CVQA Mogrovejo et al. allows using their QA data for research purposes, which is the aim of this work. We do not include the images in our release, and instead include their ID in CVQA. Refer to Mogrovejo et al. for the licenses of the images, as each has a specific license.
@@ Appendix F License
Appendix G CaMMT Data Curation Guideline
@@
Appendix H Human Preference Evaluation Instructions
@@ Appendix I CSI Retention Evaluation
    CSI Extraction: Automatically identifying culturally specific items using the prompt shown in Box I, which compares conserved translations (containing the CSI) against substituted translations (where the CSI is replaced with a more general term).
@@
    CSI Preservation Evaluation: Determining which of two competing translation systems better preserves the identified CSI when compared to a gold reference, following the evaluation setup in Box I.
For both CSI extraction and evaluation, we utilized GPT-4o with $temperature=0.0$ to ensure deterministic outputs. The CSI extraction was limited to $max\_tokens=50$, while we used default token limits for the evaluation task. All processing was performed through the OpenAI API, maintaining consistent parameters across all language pairs and translation systems.

### v2

@@ 7 Conclusions
We present CaMMT, a human-curated dataset for Multimodal Machine Translation that encompasses 19 languages across 23 regions. We evaluated five VLMs at different scales on CaMMT and observed that providing images as auxiliary context generally improves translation quality in ways that native speakers find meaningful. When translations incorporate images, they tend to better preserve cultural elements, use correct gender marking, and resolve ambiguities. All of these improvements are often overlooked by automatic MT metrics. However, we also observe a non-trivial number of cases where visual input negatively affects translations. Understanding when and why this occurs remains an important direction for future research.
@@
Appendix A Appendix
A.1 CaMMT Statistics
@@
Language-Region Script(s) Size CSI
Culturally Relevant
(non-CSI)
Not culturally
relevant
@@ A.2 Experimental Setting
A.3 Aya-Vision Evaluations
In Figure 6, we report chrF++ scores for Aya-Vision 8B and 32B. The results show a pattern consistent with Gemini, Gemma 3, and Qwen2.5-VL. In the $X\to$ En direction, providing the image generally improves performance, yielding higher scores in most cases. In the En $\to X$ direction, the effect is more mixed, though it remains mostly positive, particularly for the larger model. Overall, these results complement and reinforce our main findings.
Figure 6: Heatmaps showing average chrF++ scores in the Aya-Vision family for text+image (T+I) and text-only (T) settings. We observe a consistent behavior with the models evaluated in our main analysis.
@@ A.3 Aya-Vision Evaluations
A.4 BLEU and BertScore metrics across models
@@ A.5 Translation Prompts
A.6 Comparison between categories measured by chrF scores
@@
A.7 License
CVQA Romero et al. (2024) allows using their QA data for research purposes, which is the aim of this work. We do not include the images in our release, and instead include their ID in CVQA. Refer to Romero et al. (2024) for the licenses of the images, as each has a specific license.
@@ A.7 License
A.8 CaMMT Data Curation Guideline
@@
A.9 Human Preference Evaluation Instructions
@@ A.10 CSI Retention Evaluation
    CSI Extraction: Automatically identifying CSI using the prompt shown in Box A.10, which compares conserved translations (containing the CSI) against substituted translations (where the CSI is replaced with a more general term).
@@
    CSI Preservation Evaluation: Determining which of two competing translation systems better preserves the identified CSI when compared to a gold reference, following the evaluation setup in Box A.10.
For both CSI extraction and evaluation, we utilized GPT-4o with $temperature=0.0$ to ensure deterministic outputs. The CSI extraction was limited to max_tokens=50max\_tokens=50, while we used default token limits for the evaluation task. All processing was performed through the OpenAI API, maintaining consistent parameters across all language pairs and translation systems.